One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

Grade:12

4 Answers

SR Roy
5 years ago
Let there be x humans and y horses.Each human and horse have one head. Each human has two legs and each horse has four legs.Then,x+y=742x+4y=196which can be solved.2x+2y=148Therefore, 2y = 48y= 24X=5024humans and 50horses.
SR Roy
5 years ago
Let there be x humans and y horses.Each human and horse have one head. Each human has two legs and each horse has four legs.Then,x+y=742x+4y=196which can be solved.2x+2y=148Therefore, 2y = 48or y= 24And, x=50So, there were 50 humans and 24 horses.
SR Roy
5 years ago
Let there be x humans and y horses.Each human and horse have one head. Each human has two legs and each horse has four legs.Then,x+y=74or 2x+4y=196which can be solved.2x+2y=148Therefore, 2y = 48or y= 24And, x=50So, there were 50 humans and 24 horses.
Narendra
5 years ago

 Decomposition of H2O2 follows a first order reaction. In fifty minutes the concentration of H2O2 decreases from 0.5 to 0.125 M in one such decomposition. When the concentration of H2O2 reaches 0.05 M, the rate of formation of O2 will be